OpenStudy (anonymous):

1. Solve the absolute value equation. Check your solutions. | 5x + 13| = –7

OpenStudy (anonymous):

not likely

OpenStudy (anonymous):

not much chance that the absolute value of anything is negative, so don't even try it

OpenStudy (anonymous):

@satellite73 i need your help plz help me

OpenStudy (anonymous):

Make two cases: one where 5x+13≥0 and one where 5x+13<0 and solve each of those cases independently.

OpenStudy (anonymous):

There will be two separate equations to solve from: First when the 5x+13 is positive which means x>-13/5 which leads to +(5x+13)=-7 solving it gives the value as -4 since we 've assumed the above eqn for case x>-13/5 thus -4 is discarded. Now another eqn comes when 5x+13 is -ve means x<-13/5 which leads to -(5x+13)=-7 which gives us a solution as x=-6/5 which is discarded because previously we 'd assumed it for x<-13/5 but -6/5 is greater than -13/5

OpenStudy (anonymous):

Hence there is no solution to this equation.